i am horrible at math and i am staring blankly at this quiz question
this is the formula: ng/mL * mg / 10^6 ng * 10^3 mL / L. Figuring out the formula is as far as i got.
this is the question:
Convert
5mg/325mg to ng/ml and the answer i am looking for is the conversion to ng/ml.
Thanks i really appreciate the help.
: i am horrible at math and i am staring blankly at this quiz question
this is the formula: ng/mL * mg / 10^6 ng * 10^3 mL / L. Figuring out the formula is as far as i got.
this is the question:
Convert
5mg/325mg to ng/ml and the answer i am looking for is the conversion to ng/ml.
Thanks i really appreciate the help.

Answer by MathLover1(1157)

nano = 10^-9
milli = 10^-3
Your conversion is:
(ng/mL) *( mg / (10^6 )ng )* (10^3( mL / L))
or
(ng/mL) / 1000 = mg/L


1mg= 1000000ng
Condiser
1Kg=1L
or
1000mg=1000ml...=>...1mg=1ml
Convert
5mg/325mg to ng/ml
5000000ng /325 ml=15384.61538(ng/ml)